OpenStudy (firejay5):

Find the difference. Express your answer in simplest form. 2s/s^2-2s+1 - 2/s^2-2s+1

OpenStudy (anonymous):

Which part is the common part?

OpenStudy (firejay5):

\[\frac{ 2 }{ s^2-2s+1 } - \frac{ 2 }{ s^2 -2s+1}\]

OpenStudy (firejay5):

the first one is 2s over all of it

OpenStudy (anonymous):

Factorize the common part first s^2-2s+1 = ( s ... ) ²

OpenStudy (anonymous):

@Firejay5 questions?

OpenStudy (firejay5):

no

OpenStudy (firejay5):

I don't. I need help with this

OpenStudy (firejay5):

s-16b+1 IDK

OpenStudy (firejay5):

2s^2 -s +2

OpenStudy (firejay5):

1/2+1

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

show me, cause I don't know

OpenStudy (anonymous):

Suppose 5 candies + candies Which one is common part?

OpenStudy (firejay5):

what do you mean by common part

OpenStudy (anonymous):

Whatever the same between many elements is common part! ( in english meaning)

OpenStudy (firejay5):

candies

OpenStudy (anonymous):

Now, since candies is the common part, you factorize it out What left in the first term of 5 candies

OpenStudy (firejay5):

5 candies

OpenStudy (anonymous):

Remember the common part "candies" is take away outside, what left is inside? 5 candies + 1 candies = candies ( ......)

OpenStudy (firejay5):

6

OpenStudy (anonymous):

= candies ( 5 + 1)

OpenStudy (anonymous):

Now, back to your post, what is the common part?

OpenStudy (firejay5):

\[\frac{ 1 }{ 2+1 }\]

OpenStudy (firejay5):

I really don't know now and plus I don't think your even helping; you're making it worse for me

OpenStudy (anonymous):

What's the common part here? 2s/s^2-2s+1 - 2/s^2-2s+1

OpenStudy (anonymous):

Can u PLZ type the real question

OpenStudy (firejay5):

\[\frac{ 2s }{ s^2-2s+1 } - \frac{ 2 }{ s^2-2s+1 }\]

OpenStudy (anonymous):

@sauravshakya Thanks :)

OpenStudy (anonymous):

Now, can u take 1/(s^2 -2s +1) common

OpenStudy (firejay5):

I guess

OpenStudy (anonymous):

What do u get?

OpenStudy (firejay5):

1 - s + 1

OpenStudy (anonymous):